OpenStudy (anonymous):

30 basketballs (15 Spalding, 10 Wilson, and 5 other brand-name balls) are on a basketball court. Barry Wood closes his eyes and arbitrarily picks up a ball from the court. Determine the probability that the ball selected is A Wilson.

OpenStudy (anonymous):

should I devide 30/10

OpenStudy (anonymous):

the probability that the ball selectid is A Wilson is the total Wilson balls divided by the total balls so it's 10/30 = 1/3

OpenStudy (anonymous):

all 30/10 wilsons = 1/3